There are 20 books on algebra and calculus in our library.Prove that the greatest no. of selections each of which consists of 5 books on each topic is possible only when there are 10 books on each topic in library?

Grade:11

1 Answers

sumanth
6 years ago
No; of books of algebra=10&No; of books of calculus=10Minimum books required for selection=5eachGreatest no; of selections are10c5+10c5=2*10c52*(10*9*8*7*6)/(5*4*3*2*1)=798 is the greatest number of selections